don can drive 100 miles to work in the same amount of time that it takes him to drive 60 miles to his friends house. He drives 20 mph slower to his friends house than when he drives to work. How fast does he drive to work?
Answer by jorel555(1290)

Let s be the speed don goes when going to work. Then:
100/s=60/s-20
100s-2000=60s
40s=2000
s=50 mph going to work; 30 mph going to his friends. ☺☺☺☺